# Constrained inhomogeneous spherical equations: average-case hardness

### v1

@@ 6. Spherical functions
Therefore, if it is not true that $P_{\mathcal{A}}(n)$ is $o(n^{-d})$ for some PPT algorithm ${\mathcal{A}}$ and some $d>0$, then $P_{\mathcal{A}}(n)\geq c\cdot n^{-d}$ (for some $c>0$) satisfied for infinitely many indices $n$. By Theorem 5.2 that means that there is a PPT algorithm solving ${\mathbf{SIVP}}_{\gamma=pn^{6}}$ with overwhelming probability on every lattice for infinitely many dimensions $n$, which contradicts the assumption (6). ∎
Proposition 6.2.
@@
Thus, if we can efficiently find collisions for uniformly distributed $0/1$-spherical equations $H_{\overline{c}}\in{\mathcal{H}}_{n}$ for infinitely many indices $n$, then we can efficiently find solutions for uniformly distributed ${\mathbf{CISE}}_{\{1,2,3\}}$. Then, by Theorem 5.2, there is a PPT algorithm solving ${\mathbf{SIVP}}_{\gamma=pn^{6}}$ with overwhelming probability on every lattice for infinitely many dimensions $n$, which contradicts the assumption (6). Contradiction. ∎
@@ 7. The acyclic graph word problem: average-case hardness
The reduction described above induces a system of probability measures $\{\mu_{n}\}_{n\in{\mathbb{N}}}$ on instances of ${\mathbf{AGWP}}$ over the class of groups ${\mathbb{Z}}_{p}^{n}\rtimes{\mathbb{Z}}_{p}^{\ast}$. In fact, each $\mu_{n}$ gives uniform distribution on a subset of instances of ${\mathbf{AGWP}}$ of type shown in Figure 7.1. That establishes a one-to one correspondence between uniformly randomized ${\mathbf{CISE}}_{\{1,2\}}$ described in Section 5 and ${\mathbf{AGWP}}$ endowed with $\{\mu_{n}\}_{n\in{\mathbb{N}}}$, which implies average-case hardness of ${\mathbf{AGWP}}$ (again, assuming (6)).
@@ 8.1. Spherical equations over finite groups: constrained versus unconstrained
Problem 8.2.
Does $\{\operatorname{{GL}}(2,p)\}_{p\in\mathbb{P}}$ satisfy the second condition of Problem 8.1?
Problem 8.2 can be generalized as follows.
Problem 8.3.
Does any family of classical finite groups (such as $\operatorname{{SL}}(n,p)$, $\operatorname{{PSL}}(n,p)$, etc., see [24]) satisfy conditions of Problem 8.1?
@@ 8.3. Constrained spherical equations: generalization
As we have shown in Section 7, instances of ${\mathbf{CISE}}$ can be transformed into instances of ${\mathbf{AGWP}}$ of certain type (shown in Figure 7.1). Let us generalize the digraph shown in Figure 7.1 and modify it into a finite-state transducer. For a table of elements $C=\{c_{ij}\}$ from $G$ (where $i=1,\dots,m$ and $j=0,1$) define a transducer $\Gamma_{C}$ shown in Figure 2.
@@
Problem 8.6.
Investigate average-case hardness of computing $J_{C}^{-1}$.
The most promising approach to Problem 8.6 appears to be via self-reductions similar to those discussed in Section 8.2. Furthermore, it might be easier to design self-reductions for functions $J_{C}$ because $J_{C}$ does not require pairs $c_{i0}$ and $c_{i1}$ to be conjugate.
Problem 8.7.

### v2

@@ 6. Spherical functions
Therefore, if it is not true that $P_{\mathcal{A}}(n)$ is $o(n^{-d})$ for some PPT algorithm ${\mathcal{A}}$ and some $d>0$, then $P_{\mathcal{A}}(n)\geq c\cdot n^{-d}$ (for some $c>0$) satisfied for infinitely many indices $n$. By Theorem 12 that means that there is a PPT algorithm solving ${\mathbf{SIVP}}_{\gamma=pn^{6}}$ with overwhelming probability on every lattice for infinitely many dimensions $n$, which contradicts the assumption (6).
Proposition 14.
@@
Thus, if we can efficiently find collisions for uniformly distributed $0/1$-spherical equations $H_{\overline{c}}\in{\mathcal{H}}_{n}$ for infinitely many indices $n$, then we can efficiently find solutions for uniformly distributed ${\mathbf{CISE}}_{\{1,2,3\}}$. Then, by Theorem 12, there is a PPT algorithm solving ${\mathbf{SIVP}}_{\gamma=pn^{6}}$ with overwhelming probability on every lattice for infinitely many dimensions $n$, which contradicts the assumption (6). Contradiction.
@@ 7. The acyclic graph word problem: average-case hardness
The reduction described above induces a system of probability measures $\{\mu_{n}\}_{n\in{\mathbb{N}}}$ on instances of ${\mathbf{AGWP}}$ over the class of groups ${\mathbb{Z}}_{p}^{n}\rtimes{\mathbb{Z}}_{p}^{\ast}$. In fact, each $\mu_{n}$ gives uniform distribution on a subset of instances of ${\mathbf{AGWP}}$ of type shown in Figure 15. That establishes a one-to one correspondence between uniformly randomized ${\mathbf{CISE}}_{\{1,2\}}$ described in Section 5 and ${\mathbf{AGWP}}$ endowed with $\{\mu_{n}\}_{n\in{\mathbb{N}}}$, which implies average-case hardness of ${\mathbf{AGWP}}$ (again, assuming (6)).
@@ 8.1. Spherical equations over finite groups: constrained versus unconstrained
Problem 17.
Does $\{\operatorname{{GL}}(2,p)\}_{p\in\mathbb{P}}$ satisfy the second condition of Problem 16?
Problem 17 can be generalized as follows.
Problem 18.
Does any family of classical finite groups (such as $\operatorname{{SL}}(n,p)$, $\operatorname{{PSL}}(n,p)$, etc., see [23]) satisfy conditions of Problem 16?
@@ 8.3. Constrained spherical equations: generalization
As we have shown in Section 7, instances of ${\mathbf{CISE}}$ can be transformed into instances of ${\mathbf{AGWP}}$ of certain type (shown in Figure 15). Let us generalize the digraph shown in Figure 15 and modify it into a finite-state transducer. For a table of elements $C=\{c_{ij}\}$ from $G$ (where $i=1,\dots,m$ and $j=0,1$) define a transducer $\Gamma_{C}$ shown in Figure 2.
@@
Problem 21.
Investigate average-case hardness of computing $J_{C}^{-1}$ for $C$ sampled uniformly randomly.
The most promising approach to Problem 21 appears to be via self-reductions similar to those discussed in Section 8.2. Furthermore, it might be easier to design self-reductions for functions $J_{C}$ because $J_{C}$ does not require pairs $c_{i0}$ and $c_{i1}$ to be conjugate.
Problem 22.